# 4D Millimeter-Wave Radar in Autonomous Driving: A Survey

### v3

@@ VI-A1 Ego-Velocity Estimation
VI-A2 Dynamic Points Removal
@@
VI-A3 Angular Resolution Improving
@@
VI-A4 Network Enhancement
In the context of learning-based localization and mapping, Doppler information also holds significant value.As previously discussed, Doppler velocity can serve as an indicator of whether a point originates from a stationary or dynamic object, which enlightens authors of [83] and [95] to establish a velocity-aware attention module. This module leverages Doppler information to learn attention weights, thereby distinguishing between stationary and dynamic points.
VI-B Traditional Methods
Traditional localization and mapping refers to methods with no learning-based modules. Given the unique characteristic of 4D mmWave radars, related localization and mapping research mostly feature in odometry estimation and loop closure detection.
VI-B1 Odometry Estimation
@@
Considering the inherent noise and sparsity of 4D mmWave radar point clouds, initial odometry estimation research have primarily concentrates on the estimation of ego-velocity derived from the Doppler information instead of point cloud registration. Doer and Trommer make plenty of contributions to this field using Unmanned Aerial Vehicles (UAV). They fuse ego-velocity estimated by LSQ with the IMU data to perform UAV odometry estimation[89, 96], and further extend their work to multiple radars [97, 98] and radar-camera fusion systems [84]. As early investigations, these research efforts exhibit certain limitations. They rely on Manhattan world assumptions and consider the surroundings to be stationary, which may restrict the applicability in challenging outdoor scenarios.
@@
The ego-velocity estimated by 4D mmWave radar point clouds has been employed by various researchers, in conjunction with additional assumptions, to achieve radar-based odometry. Ng et al. [82] present a continuous-time framework that fuse the ego-velocity from multiple radars with the measurement of an IMU. The continuity of this framework facilitates closed-form expressions for optimization and makes it well-suited for asynchronous sensor fusion. Given the relatively low elevation resolution of 4D mmWave radars in contrast to the more precise Doppler information, Chen et al.[94] propose a method to detect ground points and estimate ego velocity iteratively. Furthermore, Galeote-Luque et al. [91] combine the linear ego-velocity estimated from radar point clouds with the kinematic model of the vehicle to reconstruct the 3D motion of the vehicle. The authors assert that this method ensures short-term accuracy and rapid execution, qualifying it as an effective motion prior for more sophisticated SLAM frameworks.
@@
Recent research has shifted focus from direct odometry estimation through ego-velocity, as seen in above studies, to point cloud registration akin to traditional LiDAR odometry. Michalczyk et al.[85] pioneer the realization of 3D point matching across sparse and noisy radar scans based on the classic Hungary algorithm [99]. Additionally, emerging research in 4D mmWave radar SLAM is also delving into point cloud registration techniques. Zhuang et al. [81] develop a 4D mmWave radar inertial odometry and mapping system named 4D iRIOM employing iterative Extended Kalman Filter (EKF). To mitigate the effects of sparsity, they introduce an innovative point cloud registration method between each scan and submap. This method accounts for the local geometry of points in the current scan and the corresponding $N$ nearest points in the submap, weighting the distances between them by their covariance to achieve a distribution-to-multi-distribution effect. The results of this approach are shown in Fig. 17. Moreover, they incorporate 4D iRIOM with GNSS and propose G-iRIOM [92], which further utilizes RCS value to weight the point cloud registration. Besides Zhang et al.[87, 93] apply the pose graph to construct a 4D mmWave radar SLAM system adapted from a well-known LiDAR SLAM method named hdl_graph_slam [100]. Building upon the traditional point cloud registration algorithm Generalized Iterative Closest Point (GICP)[101], they propose an adaptive probability distribution-GICP. This adaptation assigns different covariance to each point according to its uncertainty inferred from the coordinate of each point, acknowledging that points at greater distances may exhibit increased uncertainty. This design considers not only the geometric distribution of neighboring points but also the spatial variance of each point. Another 4D mmWave radar pose graph SLAM system named 4DRaSLAM is designed by Li et al [90]. This system incorporates the probability density function of each point to develop a probability-aware Normal Distributions Transform (NDT) [102] routine for scan-to-submap point cloud registration. Notably, the ego-velocity estimated from 4D mmWave radar Doppler information is utilized as a pre-integration factor in this system to replace the role of IMU.
@@
VI-B2 Loop Closure Detection
@@
VI-C Learning-based Methods
Research on learning-based 4D mmWave radar localization and mapping has predominantly focused on odometry estimation, replacing traditional point cloud registration and pose regression by deep networks.
As the originator, Lu et al. [88] extract the features from radar point clouds and IMU data by CNNs and Recurrent Neural Networks (RNNs), respectively. Subsequently, they innovatively propose a two-stage cross-modal attention mechanism to integrate these features effectively. Additionally, an RNN is utilized to capture the long-term temporal dynamics of the system. The efficacy of their approach is further enhanced when supplemented with data from RGB or depth cameras.
Recently, 4DRO-Net [83] establishes a coarse-to-fine hierarchical optimization framework that leverages a sliding window approach to iteratively estimate and refine pose estimation. With the help of a velocity-aware attention cost volume network, a feature extraction network is constructed to extract global-level and point-level features. These features are instrumental in generating initial pose estimations and subsequent corrections. The integration of these two pose constructs facilitates the derivation of an enhanced odometry estimation. Based on 4DRO-Net, Zhuo et al. present 4DRVO-Net [95], further extracts and fuses image features with 4D mmWave radar point cloud features. Fig. 18 displays the pipeline. The adaptive fusion module employs a deformable attention-based spatial cross-attention mechanism to align each 4D mmWave radar feature with corresponding image feature to achieve optimal fusion.
Figure 18: The fusion and localization pipeline of 4DRVO-Net [95]
@@ VI-D Challenge
The Doppler velocity inherent in 4D mmWave radar point clouds has been recognized and utilized in localization and mapping to realize ego-velocity estimation and dynamic object removal, etc. However, considering adequate semantic information provided by 4D mmWave radars, such as intensity, traditional 3D registration methods like ICP may be less effective. Therefore, the exploration of learning-based methods or feature-based registration could yield more effective results. Regarding mapping, the sparsity of 4D mmWave radar point clouds presents a significant challenge.A potential solution could lie in mapping the environment using 4D tensor-level data, which may enhance the mapping performance.
@@ VII Future Trends
VII-1 Point Cloud Enhancement
As the most commonly applied data format, 4D mmWave radar point clouds undergo evident inferior quality compared with other sensors —- the characteristics of radars, such as the multi-path effect, severely degrade the fidelity of point clouds. Consequently, there is an urgent need to mitigate information loss throughout the signal processing workflow, particularly by replacing CFAR with precisely engineered learning-based algorithms. Additionally, the exploration of learning-based methods DOA estimation could supplant DBF techniques, potentially leading to super-resolution in angular estimation.
VII-2 Application Algorithms Redesign
Beyond the quality enhancement of 4D mmWave radar point clouds, the post-signal processing application algorithms also warrant significant attention. Up till now, numerous application algorithms for 4D mmWave radars have been modified from their LiDAR counterparts. The unique attributes of 4D mmWave radars, such as their velocity measuring ability and adaptability in extreme environments, should be further explored by future research. For perception tasks, the fusion of multiple modalities is unquestionably a pivotal direction for future development. However, it remains to be determined whether the integration of additional sensor modalities might weaken the robustness of 4D mmWave radard in extreme weather conditions. As for 4D mmWave radar localization and mapping, the Doppler information utilizing, and sensor fusion with LiDARs and cameras offer vast expanses of potential, ripe for discovery and innovation.
VII-3 Pre-CFAR Data Utilization
@@
Regarding the distinctive data formats throughout the 4D mmWave radar signal processing workflow, such as raw ADC data, RD maps, and 4D tensors, their utilization for perception, localization and mapping tasks represents an interesting yet largely unexplored area of research. The development of learning-based models that can effectively leverage the information contained within these data formats, while maintaining satisfactory real-time performance, could potentially emerge as a focal point in future research endeavors.
VII-4 Dataset Enriching
As with all other data-driven research domains, datasets pertaining to 4D mmWave radars play a significant role in facilitating related studies. However, existing datasets that incorporate 4D mmWave radars are relatively rare. The expansion of data formats and the enrichment of scenario diversity are two critical fields waiting for further development.

### v4

@@ V-A1 Ego-Velocity Estimation
V-A2 Dynamic Points Removal
@@
V-A3 Angular Resolution Improving
@@
V-A4 Network Enhancement
In the context of learning-based SLAM, Doppler information also holds significant value.As previously discussed, Doppler velocity can serve as an indicator of whether a point originates from a stationary or dynamic object, which enlightens authors of [82] and [8] to establish a velocity-aware attention module. This module leverages Doppler information to learn attention weights, thereby distinguishing between stationary and dynamic points.
V-B Traditional Methods
Traditional SLAM refers to methods with no neural networks, and can be composed into four modules: odometry estimation, loop closure detection, global optimization and mapping. Considering the unique characteristic of 4D mmWave radars, related research mostly feature in the former two modules, and will be discussed below.
V-B1 Odometry Estimation
@@
Considering the inherent noise and sparsity of 4D mmWave radar point clouds, original odometry estimation research have primarily concentrates on the estimation of ego-velocity derived from the Doppler information instead of point cloud registration. Doer and Trommer make plenty of contributions to this field using Unmanned Aerial Vehicles (UAV). They fuse ego-velocity estimated by LSQ with the IMU data to perform UAV odometry estimation[88, 94], and further extend their work to multiple radars [95, 96] and radar-camera fusion systems [83]. As early investigations, these research efforts exhibit certain limitations. They rely on Manhattan world assumptions and consider the surroundings to be stationary, which may restrict the applicability in challenging outdoor scenarios.
@@
Additionally, the ego-velocity estimated by 4D mmWave radar Doppler information has been explored by various researchers, in conjunction with additional assumptions, to achieve radar-based odometry. Ng et al. [81] present a continuous-time framework that fuse the ego-velocity from multiple radars with the measurement of an IMU. The continuity of this framework facilitates closed-form expressions for optimization and makes it well-suited for asynchronous sensor fusion. Given the relatively low elevation resolution of 4D mmWave radars in contrast to the more precise Doppler information, Chen et al.[93] propose a method to detect ground points and estimate ego velocity iteratively. Furthermore, Galeote-Luque et al. [90] combine the linear ego-velocity estimated from radar point clouds with the kinematic model of the vehicle to reconstruct the 3D motion of the vehicle.
@@
Recent research has shifted focus from direct odometry estimation through Doppler-based ego-velocity, as seen in above studies, to point cloud registration akin to traditional LiDAR odometry. In 4D mmWave radar point cloud registration, Doppler information utilization and noise and sparsity handling are two main concerns. Michalczyk et al. [84] pioneer the realization of 3D point registration across sparse and noisy radar scans based on the classic Hungary algorithm [97]. Additionally, emerging research in 4D mmWave radar SLAM is also delving into specialized designs of point cloud registration techniques. Zhuang et al. [7] develop a 4D mmWave radar inertial odometry and mapping system named 4D iRIOM employing iterative Extended Kalman Filter (EKF). To mitigate the effects of sparsity, they introduce an innovative point cloud registration method between each scan and submap. This method accounts for the local geometry of points in the current scan and the corresponding $N$ nearest points in the submap, weighting the distances between them by their covariance to achieve a distribution-to-multi-distribution effect. The results of this approach are shown in Fig. 16. Moreover, they incorporate 4D iRIOM with GNSS and propose G-iRIOM [91], which further utilizes RCS value to weight the point cloud registration. Besides, the pose graph optimization is applied by Zhang et al.[86, 92] to construct a 4D mmWave radar SLAM system adapted from a well-known LiDAR SLAM method named hdl_graph_slam [98]. Building upon the traditional point cloud registration algorithm Generalized Iterative Closest Point (GICP)[99], they propose an adaptive probability distribution-GICP, assigning different covariance to each point according to its uncertainty inferred from the coordinate of each point, given that points at greater distances may exhibit increased uncertainty. This design considers not only the geometric distribution of neighboring points but also the spatial variance of each point. Also considering point uncertainty, Li et al [89] propose 4DRaSLAM to incorporate the probability density function of each point to develop a probability-aware Normal Distributions Transform (NDT) [100] routine for scan-to-submap point cloud registration. Notably, the ego-velocity estimated from 4D mmWave radar Doppler information is utilized as a pre-integration factor in this system to replace the role of IMU.
@@
V-B2 Loop Closure Detection
@@
V-C Learning-based Methods
Research on learning-based 4D mmWave radar SLAM has predominantly focused on odometry estimation, replacing traditional point cloud registration and pose regression by deep networks.
As the originator, Lu et al. [87] design CNNs and Recurrent Neural Networks (RNNs) to extract the features from radar point clouds and IMU data, respectively. Subsequently, they propose a two-stage cross-modal attention mechanism to achieve feature integration. An RNN is utilized additionally to capture the long-term temporal dynamics of the system.
To make full use of Doppler information, 4DRO-Net [82] establishes a velocity-aware attention cost volume network within a coarse-to-fine hierarchical optimization framework to iteratively estimate and refine pose estimation. Global-level and point-level features are extracted to generate initial pose estimations and subsequent corrections. 4DRVO-Net [8] further extracts and fuses image features with 4D mmWave radar point cloud features. Fig. 17 displays the pipeline of 4DRVO-Net. The adaptive fusion module employs a deformable attention-based spatial cross-attention mechanism to align each 4D mmWave radar feature with corresponding image feature to achieve optimal fusion.
Figure 17: The data fusion and localization pipeline of 4DRVO-Net [8]
@@ V-D Challenge
The Doppler velocity inherent in 4D mmWave radar point clouds has been recognized and utilized in SLAM to realize ego-velocity estimation and dynamic object removal, etc. However, considering adequate semantic information provided by 4D mmWave radars, such as intensity, traditional 3D registration methods like ICP may be less effective. Therefore, the exploration of learning-based methods or feature-based registration could yield more effective results. The optimal exploitation of these semantic information within the autonomous driving sphere remains a relatively open question.
Given that radar point clouds are significantly less data-heavy than their tensor counterparts, and methodologies developed for LiDARs can be adapted to 4D mmWave radar point clouds with minimal modifications, the majority of SLAM studies preferentially employ radar point clouds as their input data instead of radar tensors. But regarding mapping, the sparsity of 4D mmWave radar point clouds presents a significant challenge. A potential solution could lie in mapping and rendering the environment using 4D tensor-level data.
VI Datasets
Public datasets play an indispensable role in the advancement of 4D mmWave radar-based algorithms, as they furnish essential platforms for the development, benchmarking, and comparative analysis of diverse algorithms, thereby stimulating research in the field. This section categorizes and introduces current available datasets containing 4D mmWave radars, which are summarized in Table III.
TABLE III: 4D mmWave radar datasets
Dataset Resolution Total Frames Labeled Frames Data Formats1 Modality2 Bounding box Tracking ID Odometry
Azi. Ele. Range(m) Velo.(m/s)
Datasets for Perception
Astyx [76] N/M3 N/M N/M N/M 0.5K 0.5K RPC RCL 3D $\times$
RADIal [64] 0.1 1 0.2 0.1 25K 8.3K ADC, RAD, RPC RCL 2D $\times$
VoD [54] 1.5 1.5 0.2 0.1 8.7K 8.7K RPC RCI 3D
TJ4DRadSet [25] 1 1 0.86 N/M 40K 7.8K RPC RCL 3D
K-Radar [65] 1 1 0.46 0.06 35K 35K 4DRT, RPC RCLI 3D
Dual Radar [101] 1.2 2 0.22 N/M 50K 10K RPC RCL 3D $\times$
SCORP [77] 15 30 12 0.33 3.9K 3.9K ADC, RAD, RPC RC $\times$ $\times$
Radatron [41] 1.2 18 0.05 N/M 152K 16K RA RC 2D $\times$ $\times$
Datasets for SLAM4
Coloradar [19] 1 22.5 0.12 0.25 108K - ADC, RAE, RPC RLI $\times$ $\times$
MSC-RAD4R [2] 1 0.5 0.86 0.27 90K - RPC RCLI $\times$ $\times$
NTU4DRadLM [102] 0.5 0.1 0.86 N/M 61K - RPC RCLIT $\times$ $\times$
  • 1ADC: raw radar data after Analog-to-Digital Converter; RA: Range-Azimuth map; RD: Range-Doppler map; RAD: Range-Azimuth-Doppler cube;
    RAE: Range-Azimuth-Elevation cube; 4DRT: Range-Azimuth-Elevation-Doppler Tensor; RPC: Radar Point Cloud;
  • 2R: Radar, C:Camera, L:LiDAR, I:IMU(Inertial Measurement Unit), T:Thermal Camera
  • 3N/M: Not Mentioned.
  • 4 Datasets designed only for SLAM contain no labels like bounding box and tracking ID.
VI-A Datasets for Perception
Datasets for 4D mmWave radar perception typically include 3D (or 2D) bounding boxes for object detection tasks, and tracking ID for object tracking tasks. Astyx [76] represents the first 4D mmWave radar dataset. It consists of 500 synchronized frames (radar, LiDAR, camera) encompassing approximately 3,000 annotated 3D object annotations. As a pioneer dataset in the realm of 4D mmWave radars, the volume of data in Astyx is relatively limited.
In order to facilitate researchers in handling radar data in a more fundamental manner, the RADIal dataset [64] records raw radar data after Analog-to-Digital Converter, which serves as the foundation for generating various conventional radar representations, such as radar tensors and point clouds. Given that the raw ADC data is not interpretable by human eyes, the annotations in the RADIal dataset are presented as 2D bounding boxes in the image plane.
To advance 4D mmWave radar-based multi-class 3D road user detection, the VoD dataset [54] is collected comprising LiDAR, camera, and 4D mmWave radar data. It contains 8693 frames of data captured in complex urban traffic scenarios, and includes 123106 3D bounding box annotations of both stationary and dynamic objects and tracking IDs for each annotated object.
In a similar vein, the Tj4DRadSet dataset [25] comprises 44 consecutive sequences, totaling 7757 synchronized frames, well-labeled with 3D bounding boxes and trajectory IDs. Notably, the TJ4DRadSet dataset offers occlusion and truncation indicators in each object to distinguish between different levels of detection difficulty. Unlike the VoD dataset, Tj4DRadSet is characterized by its inclusion of a broader and more challenging array of driving scenario clips, such as urban roads, highways, and industrial parks.
To the best of our knowledge, K-Radar dataset[65] currently contains most diverse scenarios in 4D mmWave radar datasets as it collects 35k frame under a variety of weather conditions, including sunny, foggy, rainy, and snowy. K-Radar not only provides 4D mmWave radar data but also includes high-resolution LiDAR point clouds, surround RGB imagery from four stereo cameras, RTK-GPS and IMU data from the ego-vehicle. Fig. 18 illustrates the comparison of different modality sensors across different weather conditions. It is worth mentioning that K-radar is currently the only dataset that provides range-azimuth-elevation-Doppler tensors.
However, each of the datasets mentioned above contains only one type of radar, making it challenging for researchers to analyze and compare the performance of different 4D mmWave radars. The recently unveiled Dual Radar dataset[101], as illustrated in Figure 19, encompasses two distinct types of mmWave radars, the Arbe Phoenix and the ARS548 RDI radar. Dual Radar enables an investigation into the impact of different sparsity levels in radar point clouds on object detection performance, providing assistance in the selection of radar products.
Figure 19: The data visualization of Dual Radar dataset. [101]
In addition, several datasets enumerated in Table III provide radar data with relatively low angular resolution [77, 41]. These datasets, which are not detailed within the scope of this paper, provide alternative radar data characteristics that may be leveraged in different research contexts.
VI-B Datasets for SLAM
As its ability of perception in severe environments, the 4D mmWave radar enhances robust localization in difficult conditions, hence leading to the release of several 4D mmWave radar datasets specifically designed for localization, mapping and SLAM. Besides, any above dataset containing odometry information, not just those designed for SLAM, can also be employed for SLAM.
ColoRadar [19] comprising approximately 2 hours of data from radar, LiDAR, and the 6-DOF pose ground truth. It provides radar data of three processing levels: raw ADC data, 3D range-azimuth-elevation tensors derived by compressing the Doppler dimension of 4D radar tensors, and radar point clouds. This dataset collects data from a variety of unique indoor and outdoor environments, thus providing a diverse spectrum of sensor data.
Considering SLAM in severe environments, the MSC-RAD4R dataset [2] records data under a wide range of environmental conditions, with the same route yielding data from both clear and snowy weather for comparison. Additionally, MSC-RAD4R introduces artificially generated smoke environment data generated by a smoke machine, further emphasizing the robust capabilities of 4D mmWave radars.
The NTU4DRadLM dataset [102] is a recent contribution to this field captured using both robotic and vehicular platforms. Distinguished from its predecessors, NTU4DRadLM delivers an extensive array of localization-related sensor data, including the 4D mmWave radar, LiDAR, camera, IMU, GPS, and even a thermal camera. Furthermore, it encompasses a wide range of road conditions, encompassing structured, semi-structured, and unstructured roads, spanning both small-scale environments (e.g., gardens) and large-scale urban settings.
VI-C Challenge
Taking into account the datasets mentioned above, it becomes apparent that there is a lack of labeled radar ADC data required for deep radar detection. To address this deficiency, using synthetic data generation techniques based on 4D mmWave radar sensor models can be considered, though radar modeling is challenging due to multiple effects of radar signal processing like multi-path reflections and signal interference.
Moreover, the scales of current 4D mmWave radar datasets far below other famous autonomous driving datasets such as nuScenes[103] and ONCE[104]. For the evaluation of algorithmic generalizability and for facilitating comparative analysis with other sensor modalities, large scale 4D mmWave datasets are indispensable in the future.
@@ VII Future Trends
VII-A Noise and Sparsity Handling
Despite the superior resolution of 4D mmWave radars compared to traditional 3D radars, factors such as antenna design, power, and multi-path effects still lead to significant noise and sparsity issues, impacting the safety of autonomous driving applications.
VII-A1 Radar Data Generation
In recent years, the computer vision field has seen numerous studies on image super-resolution, and the related theories and models can be transferred to the mmWave radar domain for application. Learning-based methods show promise for generating higher-resolution radar data, such as replacing traditional signal processing steps like CFAR and DBF, is a promising direction. However, considering the large spectral data volume of mmWave radar (a single frame 4D radar tensor in the K-RADAR dataset [65] is around 200MB), related researches still necessitates improvements in real-time processing and addressing high bandwidth requirements for pre-CFAR data transmission. Compared to Transformer-based methods with quadratic time complexity, we believe that models using space state models [105, 106] with linear complexity will have better prospects in this field.
VII-A2 Application Algorithms Redesign
Current mmWave radar perception algorithms often construct dense BEV features using pillars or cylinders, which may not be optimal for noisy, lower-resolution mmWave radar data. We believe a more efficient form is to organize object features in a sparse query manner and aggregate sensor data features using attention mechanisms, a paradigm that has been verified in computer vision[107, 108, 109].
Additionally, the ’detection by tracking’ strategy[50] is also noteworthy. By leveraging Doppler information, a key feature distinguishing mmWave radar from LiDAR, dynamic information like scene flow can be estimated first. Temporal information can then be used for feature denoising and enhancement, followed by object detection head. We believe this paradigm has the potential to improve multi-object detection and tracking accuracy in complex dynamic autonomous driving scenarios.
VII-B Specialized Information Utilizing
Compared with LiDAR, the 4D mmWave radar is characterized by its specialized information, such as pre-CFAR data and Doppler information. Their comprehensive utilization holds great importance in the competition between 4D mmWave radar and LiDAR technologies.
VII-B1 Pre-CFAR Data
@@
Regarding the distinctive data formats throughout the 4D mmWave radar signal processing workflow before CFAR, such as raw ADC data, RD maps, and 4D tensors, their utilization for perception and SLAM tasks represents an interesting yet largely unexplored area of research. The development of learning-based models that can effectively leverage the information contained within these data formats, while maintaining satisfactory real-time performance, could potentially emerge as a focal point in future research endeavors.
VII-B2 Doppler Information
The velocity measurement ability based on Doppler effect makes the 4D mmWave radar a unique sensor. Separating point clouds from static and dynamic empowers applications such as object detection, semantic segmentation, localization, and so on. Existing research has already made certain explorations into the utilization of Doppler information, but there are still great potential of it. For example, Doppler-based multiple object tracking may achieve higher accuracy than traditional methods. Besides, Doppler information is also an indispensable feature in network designing. Related architectures such as Doppler-based attention can promote better feature extraction of the scene.
VII-C Dataset Enriching
As with all other data-driven research domains, datasets pertaining to 4D mmWave radars play a significant role in facilitating related studies. Though as Section VI illustrates, there are already several datasets dedicated in 4D mmWave radar, the scale of each dataset and the standardization of 4D mmWave radars are still two main concerns about dataset.
VII-C1 Scale Expansion
The largest 4D mmWave radar dataset contains only 152K labeled frames [41]. As a comparison, there are 1.4M and 7M frames of image in nuscenes[103] and ONCE [104] dataset, respectively. To ensure the generalizability of 4D mmWave radar algorithms, dataset scale expansion is non-negligible.
VII-C2 Standardization
The 4D mmWave radar is industry just emerging recently, with many manufacturers being newly established startups. This brings about the problem of the type standardization of 4D mmWave radars. Existing datasets contains varying types of 4D mmWave radars with different parameters such as angular resolution and the largest detection range, which hinders the cross testing in different datasets. Therefore, the type standardization of 4D mmWave radars is also of great necessity.
VII-D Tasks Exploring
Although the integration of 4D mmWave radar into autonomous driving systems has shown promising advancements, we still find certain critical applications have not yet been extensively explored.
VII-D1 Scene Reconstruction and Generation
Scene reconstruction and generation are pivotal in synthesizing realistic scenarios from actual vehicular data, allowing for object manipulation within these scenarios to generate new data sets. These techniques are instrumental in producing a substantial amount of realistic data, accelerating testing cycles, and greatly alleviate the long-tail data challenges faced by autonomous driving. In recent years, numerous algorithms for scene reconstruction and generation tasks based on Neural Radiance Field(NeRF)[110] or 3D Gaussian Splatting[111] have emerged within the visual domain, but research integrating these methods with 4D mmWave radar remains sparse. The primary challenges in adapting these visual-based methods to mmWave radar technology include their lack of sensitivity to the electromagnetic wave reflection properties of different materials, resulting in discrepancies between the synthetically rendered point clouds and those captured by actual mmWave radar sensors.
VII-D2 4D Occupancy Prediction
Occupancy Prediction is an emerging task aimed at detecting irregularly-shaped and out-of-vocabulary objects, providing detailed occupancy states and semantic information for each spatial grid. However, existing methods based on vision or LiDAR [112, 113, 114] lack consideration of target velocity, making them difficult to apply to downstream decision-making and planning processes in autonomous driving. The incorporation of the Doppler effect, a feature inherent to mmWave radar, presents a novel opportunity. By harnessing the Doppler dimension, 4D mmWave radar could simultaneously estimate occupancy, semantic attributes, and velocity of objects within spatial grids, offering a comprehensive scene discription that could significantly enhance the performance and reliability of autonomous driving systems.